OpenStudy (anonymous):

Integrate the following:

OpenStudy (anonymous):

\[∫ \sqrt{\frac{ 2 }{ 5x }} dx\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ (2/5)^\frac{ 1 }{ 2 } }{ x^\frac{ 1 }{ 2 }}\]

OpenStudy (blurbendy):

@bmelyk did you figure out the other one?

OpenStudy (anonymous):

is that the first step?

OpenStudy (anonymous):

well you can go about this many ways, but personally that's how i start

OpenStudy (anonymous):

okay so then what

OpenStudy (anonymous):

well the 2/5 is a constant, so you really only need to worry about x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[(\frac{ 2 }{ 5 })^\frac{ 1 }{ 2 }\int\limits_{}^{}(\frac{ 1 }{ x })^\frac{ 1 }{ 2}\]

OpenStudy (anonymous):

thats what i have so far.

OpenStudy (anonymous):

(1/x)^(1/2) is also just x^(-1/2)

OpenStudy (anonymous):

okie

OpenStudy (anonymous):

so then just integrate x^(-1/2)

OpenStudy (anonymous):

how do i go about that? lol, we just started this course last week

OpenStudy (anonymous):

ill just use my formula, right?

OpenStudy (anonymous):

um i dont know what you mean by formula, but anytime you integrate you are raising x by a power. So like x would become (x^2)/2.

OpenStudy (anonymous):

i got: \[\sqrt{\frac{ 2 }{ 5 }} \frac{ -2 }{ 3 } x ^{\frac{ -3 }{ 2 }} +c\]

OpenStudy (anonymous):

Close, but you took the derivative of x (kinda) x^(-1/2) would become 2x^(1/2) -1/2+1=1/2

OpenStudy (anonymous):

so its\[\sqrt{\frac{ 2 }{ 5 }} 2\sqrt{x}+c?\]

OpenStudy (anonymous):

yes i believe so

OpenStudy (anonymous):

can you help me with antoher'?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

okay... \[∫ (\sec^23x+\frac{ 1 }{ 3x }) dx\]

OpenStudy (anonymous):

alright, well \[\sec ^2x \] is a known derivative of tanx

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1/x is the derivative of natural log

OpenStudy (anonymous):

yes